# Scaling Laws For Dense Retrieval

### v1

@@ Abstract.
Scaling up neural models has yielded significant advancements in a wide array of tasks, particularly in language generation. Previous studies have found that the performance of neural models frequently adheres to predictable scaling laws, correlated with factors such as training set size and model size. This insight is invaluable, especially as large-scale experiments grow increasingly resource-intensive. Yet, such scaling law has not been fully explored in dense retrieval due to the discrete nature of retrieval metrics and complex relationships between training data and model sizes in retrieval tasks. In this study, we investigate whether the performance of dense retrieval models follows the scaling law as other neural models. We propose to use contrastive log-likelihood as the evaluation metric and conduct extensive experiments with dense retrieval models implemented with different numbers of parameters and trained with different amounts of annotated data. Results indicate that, under our settings, the performance of dense retrieval models follows a precise power-law scaling related to the model size and the number of annotations. Additionally, we examine scaling with prevalent data augmentation methods to assess the impact of annotation quality, and apply the scaling law to find the best resource allocation strategy under a budget constraint. We believe that these insights will significantly contribute to understanding the scaling effect of dense retrieval models and offer meaningful guidance for future research endeavors.
@@ 1. Introduction
The studies of scaling law in language data can be traced back to a century ago. In the 1920s, a couple of linguisticians discovered that the frequency of a word is proportional to the inverse of its rank when sorting vocabulary based on each word’s frequency in the corpus, which is widely known as the Zipf’s law (Adamic and Huberman, 2002; Newman, 2005). Later in the 1960s, Gustav Herdan found that the number of distinct words in a corpus approximately follows a function of the corpus size, which can be approximated with a power function. This is often referred to as the Heaps’s law(Lü et al., 2010). The discovery of these scaling laws in language has significantly influenced not only the research of linguistic, but also many studies in natural language processing (NLP) and information retrieval (IR). For example, Zipf’s law has provided inspiration for the development of several statistical retrieval models, and Heap’s law has served as the key principle for the estimation of inverted index, which serves as the foundation of many retrieval systems.
Recently, as the focus of language modeling has switched from statistic analysis to semantic representation learning, the studies of scaling law is also shifting from analyzing the statistical characteristics of text data to the modeling of empirical relationships between representation/model structures and downstream language tasks. For instance, pre-trained neural language models have received considerable attention due to their impressive language comprehension and generation performance. Particularly, pre-trained language models with an extreme large number of model parameters, i.e., large language models (LLM), have achieved a human-level performance in many language-related tasks. In addition to the design of model structures and training objectives, researchers have spent considerable effort investigating the relationships between parameter size and language generation quality of LLMs. Previous studies have shown that variations in the model size, data size and computation reliably of LLMs often significantly affect the converged loss of LLM pre-training and downstream task performance. They have identified a precise power law scaling relationship between model performance and such scaling factors. These scaling laws enable researchers and developers to empirically predict model performance under certain conditions without actually constructing the model. Since the training of modern LLMs demands substantial time and financial resources, such scaling laws are of great importance to both the research and applications of LLMs.
Besides LLMs, pre-trained neural language models have also served as the foundation of dense retrieval models. Due to their superior ability to capture the semantic similarity between term-based queries and documents, dense retrieval models (Guo et al., 2021; Lin et al., 2021) have become the most popular neural retrieval models in recent years. They have achieved the state-of-the-art retrieval performance in many IR tasks(Qu et al., 2020; Shao et al., 2023; Ma et al., 2023). Typical dense retrieval models use a transformer-based encoder, usually initialized with a pre-trained language model and trained with relevant query-document pairs annotated by humans, to encode both queries and documents into a shared, fixed-dimensional latent semantic space. In this space, relevant documents can be directly retrieved based on their similarity with the input queries.
Dense retrieval models and LLMs share many similarities in terms of structures and applications. First, both dense retrieval models and LLMs employ the Transformer-based architecture. Second, LLMs are optimized to maximize the probability of generating the target next token from the vocabulary based on a given context, while dense retrieval models are trained to maximize the probability of retrieving the target positive passage from the candidates based on a given query. Therefore, an obvious and important question is: whether the performance of dense retrieval models follows similar scaling laws as those observed on LLMs?
Interestingly, while some studies have indicated that larger models exhibit improved generalization capabilities in zero-shot dense retrieval tasks(Ni et al., 2021; Rosa et al., 2022), to the best of our knowledge, there isn’t any published literature explicitly investigating the scaling laws in dense retrieval models. We believe that there are two key challenges that prevent researchers from successfully uncovering such phenomenon in dense retrieval. First, the training of dense retrieval models often requires large-scale relevant annotations created by human annotators or users, which are usually expensive to collect. Unlike LLMs, whose training data can be unsupervisedly and effectively extracted from raw text data in a large scale (usually in petabytes), the annotated training data of dense retrieval models are usually limited (usually in gigabytes). Thus, the observed performance of dense retrieval models is often a twisted result affected jointly by model size and data size, making it difficult to be interpreted with a concrete function. Second, existing evaluation paradigms for dense retrieval models mostly estimate model quality with discrete functions such as NDCG and MAP. The discontinuity of these evaluation metrics makes them insensitive to model size and data size changes within some ranges, but extremely sensitive to changes that pass certain thresholds. This poses significant difficulties to the construction of a smooth and reliable function to fit the behaviors of dense retrieval models.
@@
In this paper, we focus on the investigation of scaling laws for dense retrieval models. We first propose to evaluate the quality of dense retrieval models with a contrastive perplexity metric. The idea of contrastive perplexity is inspired by the popular contrastive ranking loss and the analysis of token generation perplexity in LLMs. It measures the likelihood of retrieving a relevant document from a randomly sampled candidate set, and shares a similar structure with the training loss of dense retrieval models. The smooth nature of this metric considerably facilitates our subsequent analysis. Second, to disentangle the effects of model size and data size in dense retrieval, we conducted experiments with models implemented with different pre-trained language models with non-embedding parameter sizes ranging from 0.5 to 87 million, on two of the largest web search datasets, i.e., MSMARCO and T2Ranking. Experimental results show that, under proper experimental conditions, the performance of dense retrieval models follows a precise power-law scaling with respect to training factors. Figure 1 illustrates such power-law scaling with model size. To investigate the effect of annotation quality, we adopted several LLMs and weak supervision methods to generate training data for dense retrieval models. Our results indicate that the observed scaling laws of dense retrieval are uniformly valid across models trained with different types of annotation data. Additionally, we show that the joint effect of model and data sizes can be nicely fitted and predicted with a single function within a certain range. Such functions can be used to find the best resource allocation strategy given a restricted budget, and could potentially provide important insights for the practical implementation of dense retrieval models and green IR (Strubell et al., 2019)
The contributions of this paper can be summarized as the following:
We investigate and identify several scaling laws for dense retrieval models.
We conduct extensive experiments with different model sizes, data sizes, and types of annotation. Experiment results show that the scaling law of dense retrieval models is valid across a wide range of experimental settings.
We derive empirical functions to fit the relationship between model performance and multiple factors, and showcase their potential applications in several tasks.
@@ 2.1. Scaling Laws in Linguistic Language Data
Zipf’s law(Adamic and Huberman, 2002; Newman, 2005) is one of the best pieces of evidence about the existence of universal power laws in cognitive science and the social sciences. This law finds significant application in linguistics, asserting an inverse correlation between the frequency of a word’s occurrence in natural language and its rank in the frequency distribution.
Furthermore, Zipf’s law is intrinsically connected to other statistical scaling laws in linguistics, notably Heaps’ law(Lü et al., 2010; Gelbukh and Sidorov, 2001; Lansey and Bukiet, 2009). Heaps’ law delineates a sublinear growth trajectory between a text’s vocabulary size and its total word count. As the total word count increases, the rate of introducing new words diminishes, leading to a plateau in vocabulary expansion. This phenomenon is particularly significant in information retrieval, which serves as the key principle for the estimation of inverted index.
@@ 2.2. Neural Scaling Law
Neural scaling laws articulate the relationship between model size, dataset size, computational budget, and performance in neural network training. This concept, first introduced by Hestness et al.(Hestness et al., 2017), identified a power-law relationship, which was subsequently expanded upon for larger models by Kaplan et al.(Kaplan et al., 2020). Researchers further refined this concept by developing a unified formula for scaling laws, incorporating data-dependent scaling terms for compute-optimal training.(Hoffmann et al., 2022)
These empirical scaling laws offer crucial insights for training large Transformer-based models, particularly in accurately predicting losses under specific settings. Notably, results from experiments with smaller models can be extrapolated to larger ones and may even be applicable to other downstream tasks. Recent studies have diversified the application of these laws, examining different Transformer parameterizations. For instance, Clark et al.(Clark et al., 2022) explored their application in Mixture of Experts (MoE) models. Gao’s research(Gao et al., 2023) delved into the scaling effects in model optimization with Reinforcement Learning, pivotal in AI alignment.
Beyond language-centric tasks, these scaling principles have been adapted for domain-specific applications, such as speech recognition(Radford et al., 2023), computer vision(Zhai et al., 2022; Dehghani et al., 2023), and multi-modal language-vision settings(Jia et al., 2021; Pham et al., 2023; Radford et al., 2021). In Information Retrieval (IR), Ardalani et al.(Ardalani et al., 2022) investigated the application of scaling laws in Click-Through Rate (CTR) recommendation tasks, and Zhang et al.(Zhang et al., 2023) addressed their relevance in conventional ID-based sequential recommendation models. Nonetheless, there has been limited research into whether scaling laws remain applicable in dense retrieval.
@@ 2.3. Dense Retrieval
We now briefly revisit prior studies in the field of dense retrieval. The training data for dense retrieval tasks typically comprises annotated pairs, each consisting of a query and a human-labeled relevant passage. Early research primarily concentrated on effective negative sampling strategies used for dense retrieval training, such as employing random passages or the top irrelevant passages retrieved by BM25 as negative samples. ANCE(Xiong et al., 2020) introduced the concept of self-mined negatives, wherein negatives are refreshed every epoch, which greatly improved the retrieval performance, especially in the top ranking results. Furthermore, Zhan et al.(Zhan et al., 2021a, b) proposed dynamic hard negatives to further enhance both training efficiency and retrieval effectiveness. RocketQA(Qu et al., 2020) and TAS-B(Hofstätter et al., 2021) introduced knowledge distillation, utilizing a well-trained cross-encoder model to generate soft labels for training pairs.
Subsequently, researchers’ attention shifted to retrieval-oriented second-stage pre-training for the language model. Previous pre-training tasks, such as Masked Language Modeling (MLM) or Causal Language Modeling, were not specifically tailored for Information Retrieval (IR) scenarios. Therefore, dense retrieval models might be not able to fully utilize the language comprehension capacity developed through large-scale pre-training. Condenser(Gao and Callan, 2021a) and coCondenser(Gao and Callan, 2021b) employed the Sequence Contrastive Learning (SCL) task to improve the representational capability of the [CLS] token, enabling it to encapsulate a richer information content. Furthermore, RetroMAE(Liu and Shao, 2022) introduces an encoder-decoder architecture, wherein a shallow decoder compels the encoder to produce higher-quality representations. Contriever (Izacard et al., 2021) pre-trains dense retrieval models on crawled large-scale web pages using the Inverse Cloze Task (ICT) and the Independent Cropping Task, not only enhancing downstream retrieval performance with supervised fine-tuning but also demonstrating significant potential in zero-shot dense retrieval.
@@
Additionally, researchers have observed that the single, fixed-dimension vector representation in dense retrieval could become a limitation for further advancements. Various studies have explored more complex scoring techniques. ME-BERT(Luan et al., 2021), for instance, introduces multi-vector representations to enable more precise retrieval of long documents, while ColBERT(Khattab and Zaharia, 2020; Santhanam et al., 2021), on the other hand, investigates token-level vector representations and aggregates scores using a late-interaction mechanism. Other research endeavors have attempted to expand the vector dimension to vocabulary size. This expansion allows dense retrieval models to directly generate term weights, facilitating retrieval similar to sparse models.
@@
From the perspective of training data, dense retrieval often relies on query generation techniques for data augmentation. Query generation involves generating multiple relevant queries for a given passage. The most basic approach employs unsupervised heuristic methods, such as the previously mentioned Sequence Contrastive Learning (SCL) or Inverse Cloze Task (ICT). However, the quality of the weak supervision data generated by these methods is relatively low. Therefore, they are primarily used in the unsupervised pre-training phase due to their accessibility. More advanced methods leverage pre-trained language models like T5 to generate more precise relevant queries for data augmentation(Nogueira et al., 2019). Nevertheless, these generated queries are often used for document expansion to enhance the retrieval performance in lexical matching models. As training data, these queries are usually exploited in scenarios where human annotations are scarce, such as in out-of-domain situations. Recently, approaches like HyDE(Gao et al., 2022) have leveraged the exceptional language generation capabilities of Large Language Models (LLMs) for data augmentation. This generated data is of superior quality and can be used to replace the original queries or passages in retrieval tasks.
To summarize the prior explorations of dense retrieval models, we conclude that they mainly focus on techniques of enhancing retrieval performance within certain resource constraints. This includes the development of sophisticated training strategies and model architectures with fixed model sizes, as well as the incorporation of various augmented data into the dense retrieval (DR) lifecycle when high-quality human annotations are limited. However, our study aims to explore the scaling laws of dense retrieval, a direction orthogonal to previous research. Specifically, given the same training strategies, we examine how model performance varies under different scales of available resources (such as model size and the number of annotated training pairs).
Specifically, in this study, we aim to thoroughly investigate the following three research questions:
How does the variation of model sizes impact dense retrieval performance?
How does the change of annotated training data sizes influence dense retrieval performance?
@@ 3. Methedology
In this section, we first introduce the model architecture and datasets used for exploring the scaling effect of dense retrieval. We further discuss the training strategy used in the experiments and the performance evaluation metrics adopted.
@@ 3.1. Problem Formulation
We first give a formal description of the dense retrieval task. In a fixed corpus , the goal of retrieval is to identify the top relevant passages for a given query . Dense retrieval models accomplish this by employing an encoder that maps both queries and candidate passages into a shared dense embedding space. Subsequently, a scoring function, such as inner product or cosine similarity, is applied to the encoded dense vectors to model relevance:
where and denote the query and the passage, respectively. is the mapping function of the dense retrieval model parameterized by . Note that in this paper, we only focus on the dense retrieval models with the above structure as it is one of the most popular and representative structures in practice. We leave the studies of other types of dense retrieval models to future studies.
The training data for dense retrieval typically comprises a set of training queries and associated human annotations. Each query is annotated with one or more relevant passages, and the remaining unannotated passages are generally presumed irrelevant. In this paper, we adhere to this annotation standard and consider each query-positive-passage pair as an individual data point. Formally, the training set consists of data points:
where and denote the -th query in the training set and its corresponding annotated positive passage.

### v2

@@ Abstract.
Scaling laws have been observed in a wide range of tasks, particularly in language generation. Previous studies have found that the performance of large language models adheres to predictable patterns with respect to the size of models and datasets. This helps us design training strategies effectively and efficiently, especially as large-scale training becomes increasingly resource-intensive. Yet, in dense retrieval, such scaling law has not been fully explored. In this study, we investigate how scaling affects the performance of dense retrieval models. We implement dense retrieval models with different numbers of parameters, and train them with various amounts of annotated data. We propose to use the contrastive entropy as the evaluation metric, which is continuous compared with discrete ranking metrics and thus can accurately reflect model performance. Results indicate that the performance of dense retrieval models follows a precise power-law scaling related to the model size and the number of annotations across different datasets and annotation methods. Additionally, we show that the scaling laws help optimize the training process, such as resolving the resource allocation problem under a budget constraint. We believe that these findings significantly contribute to understanding the scaling effect of dense retrieval models and offer meaningful guidance for future research.
@@ 1. Introduction
The studies of scaling laws in language data can be traced back to a century ago. In the 1920s, a couple of linguisticians discovered that the frequency of a word is proportional to the inverse of its rank when sorting vocabulary based on each word’s frequency in the corpus, which is widely known as the Zipf’s law (Adamic and Huberman, 2002; Newman, 2005). Later in the 1960s, Gustav Herdan found that the number of distinct words in a corpus approximately follows a function of the corpus size, which can be approximated with a power function. This is often referred to as the Heaps’s law (Lü et al., 2010). These foundational discoveries in scaling laws have profoundly influenced research in linguistics and information retrieval. For example, Zipf’s law has inspired the development of several statistical retrieval models, and Heap’s law has served as the key principle for the estimation of inverted index, the foundation of many retrieval systems.
Recently, as language modeling has evolved from statistical analysis to the learning of semantic representations, the focus of scaling law research has also shifted from analyzing text statistics toward the training dynamics of large language models (LLMs). Significant research effort has been dedicated to examining how different factors influence model performance, such as model size, data volume, and computational capacity (Kaplan et al., 2020). Findings from these studies reveal precise power law relationships between model performance and scaling factors, which enable researchers and developers to empirically predict model performance without actually constructing the models (Achiam et al., 2023). Since the training of modern LLMs demands substantial time and financial resources, such scaling laws are of great importance in practice.
Similar to language modeling, Dense Retrieval models have emerged as a significant milestone in this transition from statistical analysis to semantic representation learning in Information Retrieval (Cai et al., 2021; Lin et al., 2021). In contrast to traditional statistical retrieval methods such as BM25 (Robertson and Walker, 1994), Dense Retrieval models are initialized with pre-trained language models and finetuned on annotation data in an end-to-end manner. They capture the semantic similarity between queries and documents, and demonstrate superior performance over traditional methods (Qu et al., 2021; Shao et al., 2023; Ma et al., 2023). However, researchers find that the effectiveness of dense retrieval models is sensitive to multiple training factors (Xiong et al., 2021; Zhan et al., 2021). Therefore, the construction of effective dense retrieval models under practical constraints (such as budget and latency requirements) is not straightforward, and more insights on the optimization process of Dense Retrieval are needed.
@@
In this paper, we investigate the scaling laws for dense retrieval models111Code is open-sourced at https://github.com/jingtaozhan/DRScale.. While some studies have indicated that larger models exhibit improved generalization capabilities in zero-shot dense retrieval tasks (Ni et al., 2022; Rosa et al., 2022), to the best of our knowledge, there isn’t any published literature explicitly discover scaling laws in dense retrieval models. Specifically, there are two challenges; (1) traditional performance metrics in retrieval tasks (e.g., NDCG) are discrete functions, which limits their ability to stably and smoothly reflect the change of model performance in practice; (2) the training process of Dense Retrieval involve multiple interrelated factors such as model size, annotation size, and annotation quality, which makes it difficult to isolate the effect of each factor separately. To this end, we first propose to evaluate the quality of dense retrieval models with a contrastive entropy metric. The idea is inspired by the popular contrastive ranking loss and the analysis of token generation perplexity in LLMs. It measures the likelihood of retrieving a relevant document from a randomly sampled candidate set, and shares a similar structure with the training loss of dense retrieval models. The smooth nature of this metric considerably facilitates our subsequent analysis. Second, to disentangle the effects of model size and data size in dense retrieval, we conducted experiments with models implemented with different pre-trained language models with non-embedding parameter sizes ranging from 0.5 to 87 million, on two of the largest web search datasets, i.e., MSMARCO and T2Ranking. Experimental results show that, under proper experimental conditions, the performance of dense retrieval models follows a precise power-law scaling with respect to training factors. Figure 1 illustrates such power-law scaling with model size. To investigate the effect of annotation quality, we adopted several LLMs and weak supervision methods to generate training data for dense retrieval models. Our results indicate that the observed scaling laws of dense retrieval are uniformly valid across models trained with different types of annotation data. Additionally, we show that the joint effect of model and data sizes can be nicely fitted and predicted with a single function within a certain range. Such functions can be used to find the best resource allocation strategy given a restricted budget, and could potentially provide important insights for the practical implementation of dense retrieval models and green IR (Strubell et al., 2019).
@@ 2. Background and Related Work
In this section, we revisit the background about scaling laws and dense retrieval. We start with the scaling laws in linguistic analysis and in neural language models. Then we present the explorations about dense retrieval techniques and its training technique.
@@ 2.1. Scaling Laws in Linguistic Language Data
Zipf’s law (Adamic and Huberman, 2002; Newman, 2005) is a well-known evidence about the existence of universal power laws in cognitive science and the social sciences. It shows an inverse correlation between the frequency of a word’s occurrence in natural language and its rank in the frequency distribution. It is widely applied in different areas. Furthermore, Zipf’s law is tightly connected to other statistical scaling laws in linguistics, notably Heaps’ law (Lü et al., 2010; Gelbukh and Sidorov, 2001; Lansey and Bukiet, 2009). Heaps’ law shows a sublinear growth trajectory between a text’s vocabulary size and its total word count. As the total word count increases, the rate of introducing new words diminishes, leading to a plateau in vocabulary expansion. This phenomenon is particularly significant in information retrieval, which serves as the key principle for the estimation of inverted index.
@@ 2.2. Neural Scaling Law
Neural scaling law describes the relationship between model size, dataset size, computational budget, and performance in neural network training. This concept was first introduced by Hestness et al. (2017) as a power-law relationship. Subsequently, Kaplan et al. (2020) expanded it to larger models. Hoffmann et al. (2022) further refined it by developing a unified formula for scaling laws, incorporating data-dependent scaling terms for compute-optimal training.
These empirical scaling laws offer crucial insights for training large Transformer-based models, particularly by accurately predicting loss. Notably, experimental results from smaller models can be extrapolated to larger ones. Recent studies show that such scaling laws also hold for many other model architectures. For instance, Clark et al. (2022) investigated the scaling laws in Mixture of Experts (MoE) models. Gao et al. (2023) showed the scaling effects in model optimization with Reinforcement Learning.
Beyond language-centric tasks, these scaling principles have been adapted for domain-specific applications, such as speech recognition (Radford et al., 2023), computer vision (Zhai et al., 2022; Dehghani et al., 2023), and multi-modal language-vision settings (Jia et al., 2021; Pham et al., 2023; Radford et al., 2021). In Information Retrieval (IR), Ardalani et al. (2022) investigated the application of scaling laws in Click-Through Rate (CTR) recommendation tasks, and Zhang et al. (2023) addressed their relevance in conventional ID-based sequential recommendation models. Nonetheless, there has been limited research into whether scaling laws remain applicable in dense retrieval.
@@ 2.3. Dense Retrieval
We now briefly revisit prior studies in the field of dense retrieval. The training data for dense retrieval tasks typically comprises annotated pairs, each consisting of a query and a human-labeled relevant passage. Early research primarily concentrated on effective negative sampling strategies used for dense retrieval training, such as employing random passages or the top irrelevant passages retrieved by BM25 as negative samples (Karpukhin et al., 2020). ANCE (Xiong et al., 2021) utilized self-mined hard negatives and substantially improved the retrieval performance. Furthermore, Zhan et al. (2021) proposed dynamic hard negatives to further enhance both training efficiency and retrieval effectiveness. RocketQA (Qu et al., 2021) and TAS-B (Hofstätter et al., 2021) introduced knowledge distillation, utilizing a well-trained cross-encoder model to generate soft labels for training pairs.
@@
Beyond the design of finetuning methods, researchers also explore other techniques, such as pretraining methods and multi-vector retrieval. (1) Pretraining studies design objectives that are similar to the retrieval tasks. For example, Condenser (Gao and Callan, 2021) and coCondenser (Gao and Callan, 2022) use the Sequence Contrastive Learning task to improve the representational capability. RetroMAE (Xiao et al., 2022) leverages an encoder-decoder architecture, wherein a shallow decoder encourages the encoder to produce higher-quality representations. Contriever (Izacard et al., 2022) pre-trains dense retrieval models with Inverse Cloze Task and the Independent Cropping Task. (2) Since the single vector representation in dense retrieval could become a limitation, various studies have explored more complex scoring techniques. ME-BERT (Luan et al., 2021) introduces multi-vector representations to enable more precise retrieval of long documents. ColBERT (Khattab and Zaharia, 2020; Santhanam et al., 2022) investigates token-level vector representations and aggregates scores using a late-interaction mechanism. Other researchers attempt to expand the vector dimension to vocabulary size (Formal et al., 2021b, a). This expansion allows dense retrieval models to directly generate term weights, facilitating retrieval similar to sparse models.
Prior explorations of dense retrieval models mainly focus on techniques with a static setup, such as a certain model size, certain data size, etc. Instead, we employ a dynamic setup and explore how model perform when the model size and data size are scaled.
2.4. Query Generation
@@
Besides human-labeled data, dense retrieval can also utilize query generation techniques to generate pseudo annotations (Wang et al., 2022; Ma et al., 2021). Query generation involves generating multiple relevant queries for a given passage (Nogueira et al., 2019a, b). The most basic approach employs unsupervised heuristic methods, such as the previously mentioned Sequence Contrastive Learning (SCL) or Inverse Cloze Task (ICT). However, the quality of the weak supervision data generated by these methods is relatively low. Therefore, they are primarily used in the unsupervised pre-training phase due to their accessibility. More advanced methods leverage pre-trained language models like T5 to generate more precise relevant queries for data augmentation (Nogueira et al., 2019a). Nevertheless, these generated queries are often used for document expansion to enhance the retrieval performance in lexical matching models. As training data, these queries are usually exploited in scenarios where human annotations are scarce, such as in out-of-domain situations.
@@ 3. Methedology
In this section, we first introduce the model architecture and datasets used for exploring the scaling effect of dense retrieval. We further discuss the training strategy used in the experiments and the proposed performance evaluation metrics.
@@ 3.1. Problem Formulation
We first formalize the dense retrieval model. For a given corpus, the goal is to identify the top relevant passages for a specific query. Dense retrieval models accomplish this by employing an encoder that maps both queries and candidate passages into a shared dense embedding space. Subsequently, a scoring function, such as inner product or cosine similarity, is applied to the encoded dense vectors to compute relevance scores. Let and be the query and the passage, respectively. We use to denote the the mapping function of the dense retrieval model parameterized by . The relevance score is as follows:
In this paper, we only consider that the encoders for queries and passages are shared, as it is a popular implementation choice in practice. We leave the studies of separate query and document encoders to future studies.
The training data for dense retrieval typically comprises a set of training queries and associated human annotations. Each query is annotated with one or more relevant passages, and the remaining unannotated passages are generally presumed irrelevant. In this paper, we adhere to this annotation standard and consider each query-positive-passage pair as an individual data point. Formally, the training set consists of data points, , where and denote the -th query in the training set and its corresponding annotated positive passage.